# Standard Gaussian Process Can Be Excellent for High-Dimensional Bayesian Optimization

### v4

@@ 5 Robust Initialization
Alternative method. Recent work [13] proposes an intuitive approach to address the learning challenge of the SE kernel in the high-dimensional spaces by constructing a log-normal prior over each length scale, $p(\ell_{k})=\text{LogNormal}(\ell_{k}|\mu_{0}+\frac{\log(d)}{2},\sigma_{0})$. This prior regularizes the length-scale to be on the order of $\sqrt{d}$ during training. Using our theoretical framework, this method can be justified as an alternative strategy to alleviate the gradient vanishing issue. However, our method, directly motivated by our analysis, is even simpler — requiring no additional prior construction or associated parameters, and has also shown effective in addressing the gradient vanishing issue. The comparison with  [13] is given in Section 6.
@@ 6.1 Experimental Settings
For a comprehensive evaluatio of the standard BO, we employed eleven benchmarks, of which six are synthetic and five are real-world problems. For clarity, all the tasks aim for function maximization.
$\text{Fun}(d,d^{\prime})$ Structure
Ackley(300, 150) Partial variable dependency
Rosenbrok(300, 100) Partial variable dependency
Nonoverlap additive decomposition
Hartmann6(300,6) Partial variable dependency
Rosenbrock(100, 100) Nonoverlap additive decomposition
Stybtang(200, 200)
Ackley(150, 150) None
Synthetic Benchmarks. We considered four popular synthetic functions: Ackley, Rosenbrock, Hartmann6, and Stybtang. Definitions of these functions are provided in Section B.1 of the Appendix. Each task is represented by “$\text{Fun}(d,d^{\prime})$”, where $d$ is the input dimension that BO methods optimize for, and $d^{\prime}$ is the number of effective variables ($d^{\prime}\leq d$). The ground-truth of the target function values is computed with the first $d^{\prime}$ variables. The benchmarks and the structures within the target functions are summarized in Table 3.
Real-World Benchmarks. We employed the following real-world benchmark problems: Mopta08 (124)  [14], a vehicle design problem. The objective is to minimize the mass of a vehicle with respect to 124 design variables, subject to 68 performance constraints.
We followed [6] to encode the constraints as a soft penalty, which is added into the objective. SVM (388) [6], a hyper-parameter tuning problem that optimizes 3 regularization parameters and 385 length-scale parameters for a kernel support vector machine regression model. Rover (60), a rover trajectory planning problem from [30]. The goal is to find an optimal trajectory determined by the location of 30 waypoints in a 2D environment. DNA (180) [26], a hyper-parameter optimization problem that optimizes 180 regularization parameters for weighted lasso regression on an DNA dataset [20]. Prior analysis [26] shows that only around 43 variables are relevant to the target function. NAS201 (30)  [5], a neural architecture search problem on CIFAR-100 dataset.
Methods. We tested with three versions of standard BO:
@@
  • SBO-SE(RI): GP regression with ARD SE kernel, using our proposed robust initialization $\ell_{0}=c\sqrt{d}$ where $d$ is the input dimension and $c=1$
We used UCB as the acquisition function where the exploration level $\lambda$ was set to 1.5. We used GPyTorch for GP training. To expose the possible gradient vanishing issue , we used Uniform(0.001, 30) over each length-scale. The noise variance was assigned a diffused Gamma prior, Gamma(1.1, 0.05), and the amplitude the prior Gamma(2.0, 0.15). We ran L-BFGS to obtain MAP estimation. The maximum number of iterations was 15K and tolerance level 2E-9.
We compared with six state-of-the-art high-dimensional BO methods, including Tree UCB [10], RDUCB [33], HESBO [23], ALEBO [18], SaasBO [6], and TURBO [7]. The first five have been introduced in Section 3. TURBO is designed for a special scenario where the target function can be extensively evaluated, e.g., tens of thousands evaluations. This is uncommon in BO applications, as the function evaluation is typically deemed expensive, and one aims to evaluate the target function as few as possible. TURBO searches for trust-regions in the input domain, learns a set of local GP models over each region, and then uses an implicit bandit strategy to decide which local optimization runs to continue. In addition, we tested the most recent method VBO (Vanilla BO[13] that levels up the performance of the standard BO by constructing a log-normal prior over the length-scale parameters (see Section 5). We used the original implementation and default (recommended) settings for the competing methods. Besides, we compared with two BO methods based on Bayesian neural networks (BNN). One uses Laplace approximation and the other Hamiltonian Monte-Carlo (HMC) sampling for posterior inference. We used a high-quality open-source implementation for each BNN based method. Appendix Table 6 summarizes the structural assumptions made by all the methods. The details of all the competing methods are given in Appendix Section B.2.
For each optimization task, we randomly queried the target function for 20 times to collect the initial data. We tested TURBO with a single trust region and five trust regions, denoted by TURBO-1 and TURBO-5 respectively. For HESBO and ALEBO, we varied $d_{\text{emb}}\in\{10,20\}$ and denote the choice as HESBO/ALEBO-{10, 20}. For SaasBO, we performed tests using both NUTS and MAP estimation for surrogate training, denoted by SaasBO-NUTS and SaasBO-MAP, respectively. For a reliable comparison, we ran every method for ten times.
@@ 6.2 Optimization Performance
Ultimate Outcomes. We first examined the ultimate optimization outcomes of each method. The results are reported in Fig. 3 for synthetic benchmarks and in Fig. 4 for real-world benchmarks. Due to space limit, we supplement the BNN-based BO results in Appendix Section C.1. Following [6], we used violin plots that illustrate the average of the obtained maximum function value over the ten runs (the middle bar), the best (the top bar), the worst (the bottom bar), and the distribution of the obtained maximum (the shaded region).
@@
On average, SBO-Matérn, SBO-SE(RI) and VBO consistently deliver top-tier results, often outperforming the other methods by a large margin, e.g., on benchmarks Ackley (300, 150), DNA, and SVM. It is noteworthy that SBO-SE also performed exceptionally well when the input dimension $d\leq 180$, as seen in benchmarks Rosenbrock (100, 100), Mopta08, NAS, and DNA. However, as $d$ increases, e.g., Ackley (300, 150), Hartman6 (300, 6), Rosenbrock (300, 100), StybTang(200, 200) and SVM (388), the performance of SBO-SE dropped dramatically due to the gradient vanishing issue (see our theoretical analysis and numerical experiments in Section 4). After applying our robust initialization, the performance of SE kernels, namely, SBO-SE (RI) immediately recovered to top-tier levels, closely matching both SBO-Matérn and VBO. On the other hand, though SBO-Matérn used a small length-scale initialization ($\ell_{0}=0.693$), the Matérn kernel’s structure has made it far less susceptible to the gradient vanishing issue, enabling it to handle much higher dimensions than the SE kernel. Free from this issue, SBO-Matérn consistently achieved top-tier results.
@@
Runtime Performance. Next, we examined the runtime performance. Fig. 5 reports the average maximum function value obtained at each step and the standard deviation across all runs. Since we have 16 methods for comparison, showing the standard deviation for every method makes the figure cluttered. To ensure clarify, we show the curves and standard deviation of SBO, VBO, and the best, worst, and median performed remaining baselines so far at each step.
@@
After an initial stage, SBO-Matérn, SBO-SE (RI), and VBO consistently produced superior queries, achieving larger function values. Their performance curves are generally above those of competing methods, reflecting a more efficient optimization of the target function. This trend is particularly noticeable in benchmarks such as Ackley (300, 150), Rosenbrock (300, 100), and SVM. When $d>180$, however, the performance of SBO-SE declines, with its curve positioned toward the bottom due to the gradient vanishing issue. In benchmarks such as Rosenbrock (100, 100) and DNA (180), the curve for VBO is predominantly above those of SBO-Matérn and SBO-SE (RI), indicating even more efficient optimization. However, on Ackley (150, 150) and Ackley (300, 150), VBO’s curve is mostly below that of SBO-Matérn and SBO-SE (RI). For the remaining benchmarks, their curves are close and often overlap. Overall, in all cases, they tend to converge to similar function values.
@@
Additional Results. We conducted extensive additional evaluations and ablation studies to examine the impact of alternative GP training methods, exploration parameters in UCB, and other acquisition functions on SBO performance in high-dimensional settings. We found that MLE achieves optimization results comparable to the MAP training employed in our evaluation. However, using concentrated priors on the length-scale parameters (concentrated on small mean values) — such as the popular Gamma(3,6) prior that is the default choice in BOTorch — often leads to training failure and poor performance, aligning with our theoretical analysis. Both MLE and MAP perform similarly to Markov Chain Monte Carlo (MCMC) training in terms of optimization outcomes, but MCMC is significantly slower. UCB is notably robust to variations in the exploration parameter. Additionally, log-EI facilitates strong performance in high-dimensional optimization, whereas EI and TS often lead to subpar results. Detailed results and discussions are provided in Appendix Section C.
Figure 3: Optimization performance in synthetic benchmarks. The results were summarized from 10 runs.
Figure 4: Optimization performance in real-world problems. The results were summarized from 10 runs.
@@ B.1 Definition of Synthetic Functions
Ackley. We used the Ackley function defined in [29],
@@
where each $x_{i}\in[0,1]$, ${\bf A}=[A_{ij}]$ and ${\bf P}=[P_{ij}]$ are two pre-defined matrices as in [29]. The global optimum is at ${\bf x}^{\dagger}=[0.20169,0.150011,0.476874,0.275332,0.6573]$.
Figure 5: Runtime performance: Maximum Function Value Obtained vs. Number of Steps in all real-world and synthetic benchmarks. The results were averaged from 10 runs. For cleaner view, we plotted the full results only for SBO and VBO. For the remaining methods, we show best, median, and worst result at each step.
@@
B.2 Implementation
@@
  • SaasBO-MAP. The SaasBO implementation available to the public does not include the version using MAP estimation. We therefore implemented this method based on the original paper [6]. All the hyperparameter settings follow exactly the same as the original paper.
@@
B.3 Computing Environment
@@
Mopta08 (124) Rover (60) NAS201 (30) DNA (180) SVM (388)
ARD Matérn-MAP 7.39 9.99 7.81 9.77 13.55
ARD Matérn-MLE 11.23 9.84 5.76 14.34 32.16
ARD Matérn-NUTS 492.62 404.17 393.62 575.03 768.29
SaasBO-MAP 46.53 41.36 42.35 51.33 67.37
SaasBO-NUTS 985.76 375.06 229.07 1739.56 2269.50
Table 4: Average running time (in seconds) per optimization step in real-world benchmark tasks.
Rosenbrock Ackley Stybtang Rosenbrock Ackley Hartmann6
(100,100) (150, 150) (200, 200) (300, 100) (300, 150) (300, 6)
ARD Matérn-MAP 8.52 36.26 21.42 16.36 25.90 51.56
ARD Matérn-MLE 8.85 12.94 13.67 12.63 9.99 8.81
ARD Matérn-NUTS 857.21 850.23 1120.28 1276.42 1049.78 948.25
SaasBO-MAP 70.17 70.69 66.35 100.79 85.86 85.72
SaasBO-NUTS 1294.90 1621.60 1511.99 1938.53 925.80 2204.54
Table 5: Average running time (in seconds) per optimization step on synthetic benchmark tasks.
@@
Figure 7: Optimization performance using MLE training, MAP training with a uniform prior over the length-scale parameters, and NUTS on SVM (388) and DNA (180). The results were summarized from 10 runs.
Figure 8: Optimization performance using MLE training, MAP training with a uniform prior over the length-scale parameters, and NUTS on the rest of benchmarks. The results were summarized from 10 runs.
@@ C.1 BNN-based BO and Standard BO with a Single Length-Scale
Due to the space limit, we did not report the results of the BNN-based BO in the main paper. Here, we supplement the comparison with the aforementioned two BNN-based BO approaches, BNN Laplace and BNN HMC, as specified in Section B.2. In addition, we also tested SBO with a single length-scale. We used the Matérn kernel. The results are reported in Fig. 6. Across all the benchmarks, the performance of the BNN based methods is worse than the median of GP based methods. In cases, such as SVM, Hartmann6 (300, 6) and Rosenbrock, the performance of the BNN based methods is close to or overlapping with the performance of the worst GP baselines. This might be due to that under low data regimes, it is more challenging for a neural network to capture the shape of the target function and quantify the uncertainty, especially for high-dimensional targets.
@@
Interestingly, even using a single length-scale parameter, the standard BO, namely SBO-Single — as long as it does not encounter gradient vanishing — still consistently shows superior performance, which closely matches ARD Matérn across most tasks, except for Ackley(300, 150), SBO-Single struggled in finding right locations to query, leading to poor performance.
C.2 Alternative Ways to GP Training
Figure 9: Final optimization performance using a diffuse prior, Uniform(0.001, 30), vs. a concentrated Gamma prior, Gamma(3,6), over the length scale parameters.
Figure 10: Runtime optimization performance using a uniform prior, Uniform(0.001, 30), vs. a concentrated Gamma prior, Gamma(3,6), over the length scale parameters.
C.2.1 Maximum Likelihood Estimation (MLE)
In the main paper, we report the SBO performance when training GP via MAP estimation. We then examined another popular GP training approach, MLE training. We used ADAM optimization [17] [^2] We also experimented with using L-BFGS for MLE training. However, L-BFGS frequently encounters numerical issues and often fails to reliably complete model training. This is consistent with findings reported in [13]. The issue likely arises because, without constraints or regularization on parameters like the length-scales and amplitude, L-BFGS tends to explore too aggressively, leading to numerical instability. to perform MLE, with learning rate $0.1$ and the maximum number of epochs was set to $500$. We employed ARD Matérn kernel. The ultimate and running optimization performance is shown in Fig.  9 and 10. As we can see, the BO performance of using MAP and MLE training is close, and in many cases significantly overlaps, such as in DNA, SVM, Ackley (150, 150) and StybTang (200, 200).
C.2.2 MAP Training with Alternative Priors
Second, we examined MAP training with alternative prior choices over the length-scale parameters. In the main paper, we use a uniform prior, Uniform(0.001, 30). In practice, another commonly used choice is a concentrated Gamma prior, e.g., Gamma(3, 6), which is the default choice in the popular GP/BO libraries GPyTorch and BOTorch. To see how it is concentrated, in Fig. 11 we compare its probability density curve with other priors, including the aforementioned uniform prior and the diffuse Gamma priors over the amplitude and noise variance.
We observe that the prior distribution Gamma(3, 6) places a large probability mass around its mean 0.5, strongly regularizing the length-scale parameters toward this value. According to our theoretical analysis (see Section 4), this introduces a significant risk of gradient vanishing in high-dimensional settings, potentially leading to failure in GP training.
To verify the risk, we examined the alignment between the prediction and the ground-truth across the test examples from Mopta08 (124) and DNA (180). As depicted in Fig. 12 (third column), the concentrated prior leads to a severe prediction deviation from the ground-truth, indicating failure in learning the target function. This observation aligns with findings reported in the SaasBO paper [6] (see Fig.1 of that paper).
We further investigated the impact of this prior within the optimization loop. As shown in Fig. 9, after placing the prior Gamma(3,6) over each length-scale parameter, the BO performance drops substantially across many tasks, such as Mopta08 (124), SVM (388) and Ackley(150, 150), no matter using ADAM or LBFGS to conduct training.
Figure 11: Probability density curves. Gamma(2, 0.15) and Gamma(1.1, 0.5) are the priors for the amplitude and noise variance, respectively.
Figure 12: Alignment between the prediction (y-axis) and ground-truth function values (x-axis).
C.2.3 Fully Bayesian Training
Third, we investigated whether a fully Bayesian training approach would yield substantial improvements in BO. In particular, we can use MCMC to perform posterior inference of the GP parameters, and then integrate them out during the prediction. We conducted another round of BO experiments. For each step, we used NUTS to train GP with ARD Matérn. We adopted the same settings as used in SaasBO-NUTS [6]: where the mean parameter is assigned the standard Gaussian prior $\mathcal{N}(0,1)$, amplitude $\text{Gamma}(2,0.15)$, and noise variance $\text{Gamma}(0.9,10)$. We then assigned a uniform prior over each length-scale, $\text{Uniform}(0.001,30)$. The number of warm up steps was set to 512, and after the warm up stage, we ran 256 steps, and selected the samples at every 16-th step as the posterior samples. We show the running and final optimization performance in Fig 7 , and Fig. 8. We also show the average running time per optimization step in Table 4 and Table 5. It can be observed that while NUTS contributes to a slight improvement in several tasks, the enhancement is marginal.
However, running NUTS is much more computationally costly compared to doing MAP and MLE. For example, on DNA, NUTS takes 59x and 40x running time of MAP and MLE, respectively, on average in each step. Therefore, for practical purposes, using MAP or MLE for GP training appears sufficient. It is computationally efficient, and the sacrifice of optimization performance is minimal compared to much more costly training methods like NUTS.
C.3 Different Choice of $\lambda$ in UCB

### v5

@@ 5 Robust Initialization
Alternative method. Recent work (Hvarfner et al., 2024) proposes an intuitive approach to address the learning challenge of the SE kernel in the high-dimensional spaces by constructing a log-normal prior over each length scale, $p(\ell_{k})=\text{LogNormal}(\ell_{k}|\mu_{0}+\frac{\log(d)}{2},\sigma_{0})$. This prior regularizes the length-scale to be on the order of $\sqrt{d}$ during training. Using our theoretical framework, this method can be justified as an alternative strategy to alleviate the gradient vanishing issue. However, our method, directly motivated by our analysis, is even simpler — requiring no additional prior construction or associated parameters, and has also shown effective in addressing the gradient vanishing issue. The comparison with  (Hvarfner et al., 2024) is given in Section 6.
@@ 6.1 Experimental Settings
For a comprehensive evaluation of the standard BO, we employed twelve benchmarks, of which six are synthetic and six are real-world problems. For clarity, all the tasks aim for function maximization.
Synthetic Benchmarks. We considered four popular synthetic functions: Ackley, Rosenbrock, Hartmann6, and Stybtang. Definitions of these functions are provided in Section B.1 of the Appendix. Each task is represented by “$\text{Fun}(d,d^{\prime})$”, where $d$ is the input dimension that BO methods optimize for, and $d^{\prime}$ is the number of effective variables ($d^{\prime}\leq d$). The ground-truth of the target function values is computed with the first $d^{\prime}$ variables. The benchmarks and the structures within the target functions are summarized in Appendix Table 3.
Real-World Benchmarks. We employed the following real-world benchmark problems: Mopta08 (124)  (Jones, 2008), a vehicle design problem. The objective is to minimize the mass of a vehicle with respect to 124 design variables, subject to 68 performance constraints. We followed (Eriksson & Jankowiak, 2021) to encode the constraints as a soft penalty, which is added into the objective. SVM (388) (Eriksson & Jankowiak, 2021), a hyper-parameter tuning problem that optimizes 3 regularization parameters and 385 length-scale parameters for a kernel support vector machine regression model. Rover (60), a rover trajectory planning problem from (Wang et al., 2018). The goal is to find an optimal trajectory determined by the location of 30 waypoints in a 2D environment. DNA (180) (vSehić et al., 2022), a hyper-parameter optimization problem that optimizes 180 regularization parameters for weighted lasso regression on an DNA dataset (Mills, 2020). Prior analysis (vSehić et al., 2022) shows that only around 43 variables are relevant to the target function. NAS201 (30)  (Dong & Yang, 2020), a neural architecture search problem on CIFAR-100 dataset. Humanoid Standup (1003): A novel trajectory optimization benchmark based on a humanoid simulator that uses the MuJoCo physics engine (Todorov et al., 2012). The problem dimension is 1,003. The details are provided in Appendix Section B.2.
Methods. We tested with four versions of standard BO:
@@
  • SBO-Matérn (RI): GP regression with ARD Matérn-5/2 kernel, using our proposed robust initialization $\ell_{0}=c\sqrt{d}$ where $d$ is the input dimension and $c=1$.
  • SBO-SE (RI): GP regression with ARD SE kernel, using our robust initialization with $c=1$.
We used UCB as the acquisition function where the exploration level $\lambda$ was set to 1.5. We used GPyTorch for GP training and BOTorch for Bayesian optimization. To ensure efficiency, the GP was trained via point estimation of the length-scale and noise variance parameters, with the optimizer selected from L-BFGS, Adam, or RMSProp. Further details are provided in Appendix Section B.3.
We compared with six state-of-the-art high-dimensional BO methods, including Tree UCB (Han et al., 2021), RDUCB (Ziomek & Ammar, 2023), HESBO (Nayebi et al., 2019), ALEBO (Letham et al., 2020), SaasBO (Eriksson & Jankowiak, 2021), and TURBO (Eriksson et al., 2019). The first five have been introduced in Section 3. TURBO is designed for a special scenario where the target function can be extensively evaluated, e.g., tens of thousands evaluations. This is uncommon in BO applications, as the function evaluation is typically deemed expensive, and one aims to evaluate the target function as few as possible. TURBO searches for trust-regions in the input domain, learns a set of local GP models over each region, and then uses an implicit bandit strategy to decide which local optimization runs to continue. In addition, we tested the most recent method VBO (Vanilla BO(Hvarfner et al., 2024) that levels up the performance of the standard BO by constructing a log-normal prior over the length-scale parameters (see Section 5). We used the original implementation and default (recommended) settings for the competing methods. Besides, we compared with two BO methods based on Bayesian neural networks (BNN). One uses Laplace approximation and the other Hamiltonian Monte-Carlo (HMC) sampling for posterior inference. We used a high-quality open-source implementation for each BNN based method. Appendix Table 4 summarizes the structural assumptions made by all the methods. The details of all the competing methods are given in Appendix Section B.3.
For each optimization task, we randomly queried the target function 20 times to collect the initial data, except for Humanoid-Standup, where we collected 50 initial data points. We tested TURBO with a single trust region and five trust regions, denoted as TURBO-1 and TURBO-5 respectively. For HESBO and ALEBO, we varied $d_{\text{emb}}\in\{10,20\}$ and denote the choice as HESBO/ALEBO-{10, 20}. For SaasBO, we performed tests using both NUTS and MAP estimation for surrogate training, denoted by SaasBO-NUTS and SaasBO-MAP, respectively. For a reliable comparison, we ran each method ten times, ensuring that all methods used the same randomly collected initial dataset in each run (these datasets varied across different runs). Additionally, we conducted ten extra runs for all the standard BO methods and VBO to further validate their performance.
Figure 3: Optimization performance in synthetic benchmarks.
Figure 4: Optimization performance in real-world problems.
@@ 6.2 Optimization Performance
Ultimate Outcomes. We first examined the ultimate optimization outcomes of each method. The results are reported in Fig. 3 for synthetic benchmarks and in Fig. 4 for real-world benchmarks. Due to space limit, we supplement the BNN-based BO results in Appendix Section C.1. We used violin plots that illustrate the average of the obtained maximum function value over the runs (the middle bar), the best (the top bar), the worst (the bottom bar), and the distribution of the obtained maximum (the shaded region).
@@
On average, SBO-Matérn (RI), SBO-SE (RI) and VBO consistently deliver top-tier results, often outperforming the other methods by a large margin, e.g., on benchmarks Ackley (300, 150), DNA, and SVM. It is noteworthy that SBO-Matérn, despite using a small length-scale initialization ($\ell_{0}=0.693$), also achieved top-tier performance across all benchmarks except for Humanoid-Standup, where the optimization dimension exceeds 1,000. In contrast, SBO-SE failed when the input dimension $d\geq 150$ due to the gradient vanishing issue, e.g., Ackley(150, 150), Rossenbrock(300, 100), DNA, and SVM (see our theoretical analysis and numerical experiments in Section 4). The result verifies that the Matérn kernel is far less susceptible to gradient vanishing and performs much more robustly for high-dimensional problems than the SE kernel. However, when the dimension is extremely high (e.g., over 1K), the small length-scale initialization can still cause vanishing gradients and degraded optimization performance. After applying our robust initialization, both kernels achieved top-tier performance consistently, closely matching VBO.
@@
Runtime Performance. Next, we examined the runtime performance. Appendix Fig. 5 and 6 report the average maximum function value obtained at each step and the standard deviation across all runs. Since we have 16 methods for comparison, showing the standard deviation for every method makes the figure cluttered. To ensure clarify, we show the curves and standard deviation of SBO, VBO, and the best, worst, and median performed remaining baselines so far at each step.
@@
After an initial stage, SBO-Matérn (RI), SBO-SE (RI), and VBO consistently produced superior queries, achieving larger function values. So did SBO-Matérn in all the cases except Humanoid-Standup. Their performance curves are generally above those of competing methods, reflecting a more efficient optimization of the target function. This trend is particularly noticeable in benchmarks such as Ackley (300, 150), Rosenbrock (300, 100), and SVM. The performance of SBO-SE degraded when $d\geq 150$, with its curve positioned toward the bottom due to the gradient vanishing issue. In benchmarks such as Rosenbrock (100, 100) and DNA (180), the curve for VBO is predominantly above those of SBO-Matérn (RI) and SBO-SE (RI), indicating even more efficient optimization. However, on Ackley (150, 150), Ackley (300, 150) and Humanoid-Standup (1003), VBO’s curve is mostly below that of SBO-Matérn (RI) and SBO-SE (RI), and exhibits worse performance. For the remaining benchmarks, their curves are close and often overlap. Overall, in all cases, they tend to converge to similar function values.
@@
Additional Results. We conducted extensive additional evaluations to examine exploration parameters in UCB, and other acquisition functions on SBO performance in high-dimensional settings. We found that UCB is notably robust to variations in the exploration parameter. Additionally, log-EI facilitates strong performance in high-dimensional optimization, whereas EI and Thompson Sampling (TS) often lead to subpar results. Detailed results and discussions are provided in Appendix Section C.
@@ 7 Conclusion
Acknowledgement
This work has been supported by NSF CAREER Award IIS-2046295, NSF OAC-2311685, and Margolis Foundation.
@@ Appendix B More Experiment Details
$\text{Fun}(d,d^{\prime})$ Structure
Ackley(300, 150) Partial variable dependency
Rosenbrok(300, 100) Partial variable dependency
Nonoverlap additive decomposition
Hartmann6(300,6) Partial variable dependency
Rosenbrock(100, 100) Nonoverlap additive decomposition
Stybtang(200, 200)
Ackley(150, 150) None
@@ B.1 Definition of Synthetic Functions
Ackley. We used the Ackley function defined in (Surjanovic & Bingham, ),
@@
where each $x_{i}\in[0,1]$, ${\bf A}=[A_{ij}]$ and ${\bf P}=[P_{ij}]$ are two pre-defined matrices as in (Surjanovic & Bingham, ). The global optimum is at ${\bf x}^{\dagger}=[0.20169,0.150011,0.476874,0.275332,0.6573]$.
Figure 5: Runtime performance: Maximum Function Value Obtained vs. Number of Steps in all synthetic benchmarks. For cleaner view, we plotted the full results only for SBO and VBO. For the remaining methods, we show best, median, and worst result at each step.
Figure 6: Runtime performance: Maximum Function Value Obtained vs. Number of Steps in all real-world benchmarks. For cleaner view, we plotted the full results only for SBO and VBO. For the remaining methods, we show best, median, and worst result at each step.
@@
B.2 Humanoid-Standup Benchmark
We created a novel trajectory optimization benchmark in a Humanoid Standup task based on the MuJoCo physics engine (Todorov et al., 2012). The action for the environment is determined by 17 parameters (corresponding to 17 motors in the humanoid). We set the trajectory length to 59. Therefore, the dimension of the problem is $59\times 17=1,003$. The goal is to find a trajectory $\tau=({\bf a}_{1},\ldots,{\bf a}_{59})$ of motor actions that maximize the reward, for a given initial state. This is an instance of optimal control and planning, which is a classical problem in reinforcement learning. It is noteworthy that the recent work of Hvarfner et al. (2024) also created a BO benchmark (named “Humanoid") based on the MuJoCo engine, but the problem setting is different. Their benchmark seeks to optimize a linear policy, represented by a $376\times 17$ parameter matrix that linearly maps the humanoid state to an action at each step. In contrast, our benchmark makes no assumptions about an underlying policy and directly optimizes the entire trajectory, making it policy-free.
B.3 Implementation
@@
  • SaasBO-MAP. The SaasBO implementation available to the public does not include the version using MAP estimation. We therefore implemented this method based on the original paper (Eriksson & Jankowiak, 2021). All the hyperparameter settings follow exactly the same as the original paper.
@@
  • GP Training in Standard BO. For efficiency, we trained GP via point estimation. The optimizer was chosen from L-BFGS, Adam, and RMSProp. For Adam and RMSProp, we set the initial learning rate to 0.1 and 0.01, respectively. The maximum number of epochs was set to 1,500 for both methods. We maximize the log marginal likelihood of the GP regression model. The positiveness of the length-scale and noise variance parameters are ensured via SoftPlus transform. For L-BFGS, we found that maximizing the log marginal likelihood often incurs numerical issues. To achieve numerical stability, we employed the prior Uniform(0.001, 30) over each length-scale, a diffused Gamma prior over the noise variance Gamma(1.1, 0.05), and the prior Gamma(2.0, 0.15) over the amplitude. The maximum number of iterations was set to 15K and tolerance level 2E-9.
B.4 Computing Environment
@@ C.1 BNN-based BO and Standard BO with a Single Length-Scale
Due to the space limit, we did not report the results of the BNN-based BO in the main paper. Here, we supplement the comparison with the aforementioned two BNN-based BO approaches, BNN Laplace and BNN HMC, as specified in Section B.3. In addition, we also tested SBO with a single length-scale. We used the Matérn kernel. The comparison was performed across all the synthetic and real-world benchmarks except Humanoid-Standup (1003). The results are reported in Fig. 7. Across all the benchmarks, the performance of the BNN based methods is worse than the median of GP based methods. In cases, such as SVM, Hartmann6 (300, 6) and Rosenbrock, the performance of the BNN based methods is close to or overlapping with the performance of the worst GP baselines. This might be due to that under low data regimes, it is more challenging for a neural network to capture the landscape of the target function and quantify the uncertainty, especially for high-dimensional targets.
@@
Interestingly, even using a single length-scale parameter, the standard BO, namely SBO-Single — as long as it does not encounter gradient vanishing — still consistently shows superior performance, which often closely matches ARD Matérn, except for Ackley(300, 150), SBO-Single struggled in finding right locations to query, leading to poor performance.
C.2 Different Choice of $\lambda$ in UCB